OpenStudy (anonymous):

MEDAL AND FAN!! A 3-gallon cylindrical water container has a height of 1.5 feet and a radius of 3.5 inches. Nate fills up a 2-gallon jug from the container before going on a hike. If the container was full before Nate filled the jug, what is the height of the water in the container after Nate filled the jug? A. 0.5 inches B. 3 inches C. 6 inches D. 12 inches

OpenStudy (officialbluemonster):

You know I started out doing this by calculating the volume of the container (692 cu in), and then I was going to figure out how many cubic inches of water were in a gallon. But then I looked at it again. When the water goes out of the tank, the radius stays the same. Only the height changes. If there's 1.5 ft of water in a 3 gal tank, each gal is 1.5 ft / 3 gal = 0.5 ft per gallon. Nate draws out two gallons, leaving one left. The height of the water in the jug is now 0.5 ft.

OpenStudy (anonymous):

you are wrong! i tried that b4 it was wrong! @OfficialBluemonster

OpenStudy (irishboy123):

\(\Delta V = V_f - V_i\)

OpenStudy (wolf1728):

18" height and 3.5 inch radius Tank's volume would be PI * r * r * height = PI * 3.5 * 3.5 * 18 = 692.72 cubic inches. a US gallon equals 231 cubic inches https://en.wikipedia.org/wiki/Gallon Since 692.72 / 231 equals 2.99879 (practically 3) then we can say the tank is 100% full and the height would be 18 inches. I think you can solve the rest.